OpenStudy (anonymous):

convert 0.079 to a percent and fraction

OpenStudy (unklerhaukus):

\[0.079=\frac{0.079}{1}=\frac{0.79}{10}=\frac{7.9}{100}=\frac{79}{1000}\]\[\qquad\qquad\qquad\qquad\qquad\quad=7.9\%\] \[\]

OpenStudy (anonymous):

thnks alt